COULD SOMEONE SHOW ME HOW THESE WOULD BE DONE? THE QUESTION IN THE BOOK ASKS TO SIMPLIFY THESE RADICAL EXPRESSIONS. RATIONALIZE THE DENOMINATORS IF NECE SSARY A. 360 B. 3/5 I DO NOT HAVE A BUTTON FOR THE SIGN INDICATING ROOTS BUT THESE ARE IN ROOT FORM. C. 3ROOT SIGN 24/125 D. 3 ROOT SIGN 2/3 IF YOU CAN HELP I WOULD BE GRATEFULL.
Answer by stanbon(75887)

SIMPLIFY THESE RADICAL EXPRESSIONS. RATIONALIZE THE DENOMINATORS IF NECESSARY A. sqrt(360) = sqrt[36*10] = 6sqrt(10)
------------------------
B. sqrt(3/5) = sqrt(15/25) = (1/5)sqrt(15)
------------------------
C. 3ROOT SIGN (24/125) = cube root(8*3/125) = (2/5)cubert(3)
------------------------
D. 3 ROOT SIGN (2/3) = cube root (18/27) = (1/3)cubert(18)
==================
Cheers,
Stan H.